# Scene Coordinate Reconstruction: Posing of Image Collections via Incremental Learning of a Relocalizer

### v1

@@ 4.2 Mip-NeRF 360
Discussion. We present PSNR results in 3. NoPe-NeRF does not perform well on this dataset, even after processing each scene for 2 days on average. The differences between sequential images are too large. DROID-SLAM fails for the same reason. BARF performs poorly because the identity pose is not a reasonable initialisation for most scenes in this dataset. In contrast, ACE0 reconstructs the dataset successfully. While it achieves slightly lower PSNR than COLMAP, we find its pose estimates to be similar, cf. Figure 4 and the supplement.
@@ 4.3 Tanks and Temples
Frames DROID- Frames DROID- Sparse Sparse COLMAP Reality SLAM ACE0 COLMAP SLAM ACE0 COLMAP + COLMAP + (default) Capture [92]   (ours) (fast) [92]   (ours) Reloc + BA   ACE0 (ours) Training Barn 410 24.0 21.2 19.0 16.5 12.2k 24.4 13.5 17.7 26.3 25.2 Caterpillar 383 17.1 15.9 16.6 16.8 11.4k 18.6 18.9 18.6 18.7 18.8 Church 507 18.3 17.6 14.3 17.2 19.3k 12.1 11.5 16.5 18.5 18.1 Ignatius 264 20.1 17.7 17.8 19.7 7.8k 20.8 19.1 20.7 20.9 20.6 Meetingroom 371 18.6 18.1 15.6 18.0 11.1k 19.4 17.1 16.6 20.8 20.4 Truck 251 21.1 18.9 18.3 20.1 7.5k 23.6 20.6 23.0 23.4 23.3 Average 364 19.9 18.2 16.9 18.1 14.6k 19.8 16.8 18.9 21.4 21.1 Avg Time 1h 3min 5min 1.1h 74h 18min 2.2h 8h 1.8h Intermediate Family 152 19.5 18.8 17.6 19.0 4.4k 21.2 19.8 18.0 21.3 21.1 Francis 302 21.6 20.7 20.7 20.1 7.8k 19.9 21.8 21.7 22.5 22.6 Horse 151 19.2 19.0 16.3 19.4 6.0k 21.6 19.2 21.7 22.6 21.5 Lighthouse 309 16.6 16.5 13.6 17.5 8.3k 19.0 18.9 18.6 19.5 19.7 Playground 307 19.1 19.2 11.4 18.7 7.7k 17.9 11.3 20.4 21.2 20.9 Train 301 16.8 15.4 13.8 16.2 12.6k 19.6 15.6 18.5 19.8 19.2 Average 254 18.8 18.3 15.6 18.5 7.8k 19.9 17.8 19.8 21.1 20.8 Avg Time 32min 2min 3min 1.3h 48h 14min 2.2h 5h 1h Advanced Auditorium 302 19.6 12.2 16.7 18.7 13.6k 13.7 16.6 20.0 21.4 20.1 Ballroom 324 16.3 18.3 13.1 17.9 10.8k 17.2 10.4 18.9 18.0 17.3 Courtroom 301 18.2 17.2 12.3 17.1 12.6k 14.6 10.2 16.3 18.7 18.1 Palace 509 14.2 11.7 10.8 10.7 21.9k 13.8 8.6 11.0 15.3 12.1 Temple 302 18.1 15.7 11.8 9.7 17.5k 13.3 11.9 14.8 19.6 17.2 Average 348 17.3 15.0 12.9 14.8 15.6k 14.5 11.5 16.2 18.6 16.9 Avg Time 1h 2min 4min 1h 71h 27min 2.8h 10h 2.1h
@@
Next, we consider more than 1k images per scene (right side of Table 4). Here, we run COLMAP with parameters tuned for large images collections (fast) due to the extremely large image sets. ACE0 offers a reconstruction quality comparable to COLMAP on average while also being fast. RealityCapture produced fractured reconstructions when applied to these large images sets. We show some of its results in the supplement. DROID-SLAM still struggles on many scenes despite having access to sequential images that are temporally close.
@@ 5 Conclusion and Future Work
We have presented scene coordinate reconstruction, a new approach to learning-based structure-from-motion built on top of scene coordinate regression. We learn an implicit, neural scene representation from a set of unposed images. Our method, ACE0, is able to reconstruct a wide variety of scenes. In many cases, the accuracy of estimated poses is comparable to that of COLMAP. Unlike previous learning-based SfM methods, ACE0 can be applied to sets of multiple thousand unsorted images, without pose priors, and reconstructs them within a few hours.
Limitations. We show some failure cases in the supplement. Scene coordinate regression struggles with repetitive structures since the network is not able to make multi-modal predictions for visually ambiguous inputs. Scene coordinate regression struggles with representing large areas. The common solution is to use network ensembles based on pre-clustering of the scene [15, 11]. However, in a reconstruction setting, networks would have to be spawned on the fly as the reconstruction grows. In our experiments we assumed a simple pinhole camera model. To the best of our knowledge, scene coordinate regression has not been coupled with image distortion, thus far.
@@ Appendix 0.A Implementation
We start the reconstruction as explained in Section 3.3 of the main paper. We repeat neural mapping and registration of new views until all views have been registered, or less than 1% of additional views are registered compared to the last relocalization stage. In each relocalization iteration, we re-estimate the poses of all views, even if they have been registered before. This gives the pipeline the ability to correct previous outlier estimates. We consider an image to be registered successfully, if it has at least 500 inlier correspondences with an inlier threshold of 10px. At the end of the reconstruction, we apply a more conservative threshold of 1000 inliers to make sure only high quality poses are kept. For each neural mapping iteration, we load the network weights of the scene coordinate regression network of the previous mapping iteration and refine further. We conclude the reconstruction loop with a final refit of the model, using the standard ACE training parameters. In this last iteration, we start with a randomly initialised network, rather than loading previous weights.
@@
ACE uses a soft-clamped version of the reprojection error in Eq. 3 (main paper) with a soft threshold that reduces from 50px to 1px throughout training. We use the same, soft-clamped loss but keep the threshold fixed at 50px as we saw no benefit when using the loss schedule but it complicates early stopping. During the seed iteration, we switch from optimizing the Euclidean distance to optimizing the reprojection error if the reprojection error of a scene coordinate is lower than 10px. This hybrid objective is inspired by the RGB-D version of DSAC* [16].
@@ Appendix 0.B Benchmark
In the following, we give more information about the relocalization experiment of Section 4.1 and 3 of the main paper. We conducted the experiment on the 7-Scenes dataset adhering to its training / test split. Our experiments utilize the pseudo ground truth (pGT) poses of KinectFusion that come with the dataset, as well as the COLMAP pGT of [12]. The authors of [12] obtained the COLMAP pGT by reconstructing the training images of each scene with COLMAP from scratch. To make the pGT scale-metric, they aligned the COLMAP reconstruction with the KinectFusion poses. To obtain poses for the test images of each scene as well, they registered the test images to the training reconstruction. Lastly, they ran a final round of bundle adjustment to refine the test poses while keeping the training poses fixed.
@@
In our experiment, we train the ACE relocalizer on the training images of each scene. Firstly, we train ACE using the KinectFusion poses that come with the dataset. Secondly, we train ACE using the COLMAP pGT of [12]. We evaluate both version of the relocaliser using the COLMAP pGT. As expected, ACE trained with KinectFusion poses scores lower than ACE trained with COLMAP pGT, when evaluating against COLMAP pGT. This confirms the finding of [12] that KinectFusion and COLMAP produce different pose estimates.
@@
Next, we run ACE0 on the training set of each scene to reconstruct it and estimate the training poses. One output of ACE0 is an ACE relocalizer, trained from scratch without poses, in a self-supervised fashion. We apply the ACE0 relocalizer to the test images to estimate their poses - without any further model refinement or joint optimization. The ACE0 reconstruction is scale-less, or only roughly scaled based on the ZoeDepth estimate of the seed image. To enable the scale-metric evaluation of 3 of the main paper, we align the ACE0 pose estimates to the test pseudo ground truth by fitting a similarity transform. Specifically, we run RANSAC [33] with a Kabsch [45] solver on the estimated and the ground truth camera positions with 1000 iterations. Due to the optimization of the similarity transform, we might underestimate the true relocalization error. For a fair comparison, we perform the exact same evaluation procedure for our ACE baselines and fit a similarity transform although they have already been trained with scale-metric poses. The effect of fitting the similarity transform is small. ACE, trained on COLMAP pGT, achieves 97.1% average relocalization accuracy when we fit the similarity transform, and 97.6% average accuracy without fitting the transform. We compute accuracy as the percentage of test images where the pose error is smaller than 5cm and 5 compared to ground truth. The ACE0 relocalizer achieves 93.8% average accuracy when evaluated against the COLMAP pGT. The accuracy differences are smaller than 1% for all scenes except Stairs. This experiment shows that ACE0 is a viable self-supervised relocalizer, and that ACE0 poses are very close to COLMAP poses up to a similarity transform - much closer than e.g. KinectFusion poses are to COLMAP poses.
@@ Appendix 0.C Additional Results
We show qualitative results in terms of estimated poses for more scenes in Fig. 5. We show failure cases in Fig. 6 corresponding to some of the limitations discussed in Section 5 of the main paper.
@@
Additionally, we show synthesized views based on ACE0 poses. Corresponding to our quantitative results in the main paper (Table 1, 3, Table 4 of the main paper), we select one test image per scene that is representative of the view synthesis quality of ACE0 on that scene. More specifically, we select the test image where ACE0 achieves its median PSNR value. For comparison, we synthesise the same image using the estimated poses of our competitors. For synthesised images of 7-Scenes, see Fig. 7. For comparison with BARF and NoPe-NeRF, the main paper reports results on a 200 image subset per scene. We show the associated synthesized images in Fig. 8. For comparison with DUSt3R, the main paper reports results on a 50 image subset per scene. We show the associated synthesized images in Fig. 9. We show synthesized images for the MIP-NeRF 360 dataset in Fig. 10. We also show synthesized images for Tanks and Temples, in Fig. 11 for Training scenes, in Fig. 12 for Intermediate scenes, and in Fig. 13 for Advanced scenes.
@@
Figure 7: View Synthesis Quality on 7-Scenes. For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). These results correspond to Table 1 of the main paper.
@@
Figure 8: View Synthesis Quality on 7-Scenes (200 Images Subset). For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). For all methods, we obtain these results on a subset of 200 images per scene to achieve a reasonable training time of NoPe-NeRF. These results correspond to Table 1(right) of the main paper.
@@
Figure 9: View Synthesis Quality on 7-Scenes (50 Images Subset). For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). For all methods, we obtain these results on a subset of 50 images per scene to prevent DUSt3R from running out of GPU memory. These results correspond to Table 1(right) of the main paper.
@@
Figure 10: View Synthesis Quality on the MIP-NeRF 360 dataset. For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). These results correspond to 3 of the main paper.
@@
Figure 11: View Synthesis Quality on Tanks and Temples (Training Scenes). For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). These results correspond to Table 4 of the main paper. Top: Results based on 150-500 images per scenes. Bottom: Results based on 4k-22k frames sampled from the full video.
@@
Figure 12: View Synthesis Quality on Tanks and Temples (Intermediate Scenes). For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). These results correspond to Table 4 of the main paper. Top: Results based on 150-500 images per scenes. Bottom: Results based on 4k-22k frames sampled from the full video.
@@
Figure 13: View Synthesis Quality on Tanks and Temples (Advanced Scenes). For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). These results correspond to Table 4 of the main paper. Top: Results based on 150-500 images per scenes. Bottom: Results based on 4k-22k frames sampled from the full video.
@@
For completeness, we augment Table 1 in the main paper with SSIM [99] and LPIPS [106] scores for the 7-Scenes dataset, c.f., Tables 5 and 6 respectively. Similarly, Tables 14 and 15 show SSIM [99] and LPIPS [106] scores for the Tanks and Temples dataset, augmenting Table 4 in the main paper. Furthermore, Tables 17 and 17 show SSIM [99] and LPIPS [106] scores for the Mip-NeRF 360 dataset, augmenting 3 in the main paper. In all cases, SSIM and LPIPS behave very similar to PSNR in our experiments, and support the conclusions of the main paper.
Pseudo Ground Truth All Frames 200 Frames 50 Frames Frames Kinect COLMAP COLMAP DROID-SLAM ACE0 KF+ACE0 BARF NoPE-NeRF ACE0 DUSt3R ACE0 Fusion (default) (fast) [92] (ours) (ours) [53] [10] (ours) [98] (Ours) Chess 6k 0.69 0.86  0.86  0.76  0.84  0.84 0.55 0.51 0.75 0.66 0.70 Fire 4k 0.55 0.72  0.72  0.44  0.69  0.70 0.45 0.41 0.65 0.55 0.59 Heads 2k 0.69 0.82  0.82  0.78  0.81  0.81 0.54 0.52 0.77 0.69 0.79 Office 10k 0.72 0.85  0.84 failed  0.83  0.83 0.59 0.55 0.78 0.57 0.61 Pumpkin 6k 0.71 0.85  0.84  0.69  0.83  0.82 0.75 0.65 0.82 0.76 0.77 RedKitchen 12k 0.61 0.80  0.80  0.51  0.77  0.78 0.52 0.50 0.69 0.54 0.55 Stairs 3k 0.63 0.76  0.61  0.47  0.57  0.70 0.65 0.66 0.67 0.54 0.55 Average 0.66 0.81  0.79 N/A  0.76  0.78 0.58 0.54 0.73 0.62 0.65 Avg. Time realtime 38h  13h  18min  1h  7min 8.5h 47h 27min 4min* 16min
@@
Pseudo Ground Truth All Frames 200 Frames 50 Frames Frames Kinect COLMAP COLMAP DROID-SLAM ACE0 KF+ACE0 BARF NoPE-NeRF ACE0 DUSt3R ACE0 Fusion (default) (fast) [92] (ours) (ours) [53] [10] (ours) [98] (Ours) Chess 6k 0.37 0.17  0.17  0.33  0.18  0.18 0.80 0.76 0.28 0.42 0.36 Fire 4k 0.44 0.28  0.28  0.70  0.29  0.29 0.79 0.80 0.33 0.40 0.37 Heads 2k 0.42 0.26  0.26  0.32  0.26  0.27 0.75 0.70 0.31 0.40 0.28 Office 10k 0.42 0.24  0.24 failed  0.27  0.26 0.77 0.76 0.32 0.64 0.54 Pumpkin 6k 0.40 0.23  0.23  0.50  0.24  0.25 0.45 0.74 0.25 0.30 0.27 RedKitchen 12k 0.51 0.24  0.25  0.82  0.27  0.27 0.84 0.84 0.38 0.62 0.70 Stairs 3k 0.39 0.31  0.57  0.83  0.51  0.34 0.68 0.70 0.55 0.74 0.61 Average 0.42 0.25  0.29 N/A  0.29  0.26 0.73 0.76 0.34 0.50 0.45 Avg. Time realtime 38h  13h  18min  1h  7min 8.5h 47h 27min 4min* 16min
@@
We show variations of ACE0 in Table 7 alongside PSNR numbers and reconstruction times for 7-Scenes.
ACE0 picks 5 random seed images to start the reconstruction. We show average statistics of ACE0 over five runs, starting off with different sets of seed images each time. The standard deviation is low with 0.2 dB in PSNR, and about 4 minutes in terms of reconstruction time.
@@
We run a version of ACE0 that ingests ZoeDepth estimates for all frames and uses them in all reconstruction iterations. Thus, rather than optimizing the reprojection error of Eq. 3 of the main paper, we optimize the Euclidean distance to pseudo ground truth scene coordinates computed from depth estimates. Accuracy is slightly lower, and reconstruction times longer. The depth estimates are not multi-view consistent, and the model has problems to converge.
@@
Train 97.7% 100% 97.8% 100% 99.3%
Advanced Auditorium 97.7% 98.7% 98.8% 99% 98.7%
Ballroom 99.7% 100% 100% 100% 99.5%
Courtroom 93.4% 100% 94.0% 100% 98.5%
Palace 25.7% 99% 58.7% 72.3% 77.2%
Temple 6.3% 100% 77.0% 100% 96.2%

### v2

@@ 4.2 Mip-NeRF 360
Discussion. We present PSNR results in 3. NoPe-NeRF does not perform well on this dataset, despite processing each scene for 2 days. The differences between sequential images are too large. DROID-SLAM fails for the same reason. BARF performs poorly because the identity pose is not a good initialisation for most scenes. In contrast, ACE0 reconstructs the dataset successfully. While it achieves slightly lower PSNR than COLMAP, its pose estimates are similar, cf. Figure 4. The supp. shows that synthesized images based on ACE0 are visually close to those of COLMAP while our learning-based competitors are far off.
@@ 4.3 Tanks and Temples
Frames DROID- Frames DROID- Sparse Sparse COLMAP Reality SLAM ACE0 COLMAP Reality SLAM ACE0 COLMAP + COLMAP + (default) Capture [94]   (ours) (fast) Capture [94]   (ours) Reloc + BA   ACE0 (ours) Training Barn 410 24.0  21.2  19.0  16.5 12.2k 24.4  16.9  13.5  17.7  26.3  25.1 Caterpillar 383 17.1  15.9  16.6  16.9 11.4k 18.6  17.9  18.9  18.6  18.7  18.8 Church 507 18.3  17.6  14.3  17.2 19.3k 12.1  -  11.5  16.5  18.5  17.3 Ignatius 264 20.1  17.7  17.8  19.8 7.8k 20.8  18.6  19.1  20.7  20.9  20.7 Meetingroom 371 18.6  18.1  15.6  18.0 11.1k 19.4  18.2  17.1  16.6  20.8  20.3 Truck 251 21.1  19.0  18.3  20.1 7.5k 23.6  19.1  20.6  23.0  23.4  23.1 Average 364 19.9  18.2  16.9  18.1 14.6k 19.8  18.2  16.8  18.9  21.4  20.9 Avg Time 1h  3min  5min  1.1h 74h  14h  18min  2.2h  8h  1.8h Intermediate Family 152 19.5  18.8  17.6  19.0 4.4k 21.2  19.8  19.8  18.0  21.3  21.3 Francis 302 21.6  20.7  20.7  20.1 7.8k 19.9  20.4  21.8  21.7  22.5  22.7 Horse 151 19.2  19.0  16.3  19.5 6.0k 21.6  20.7  19.2  21.7  22.6  22.3 Lighthouse 309 16.6  16.5  13.6  17.5 8.3k 19.0  16.6  18.9  18.6  19.5  20.5 Playground 307 19.1  19.2  11.4  18.7 7.7k 17.9  16.5  11.3  20.4  21.2  21.0 Train 301 16.8  15.4  13.8  16.2 12.6k 19.6  14.4  15.6  18.5  19.8  18.5 Average 254 18.8  18.3  15.6  18.5 7.8k 19.9  18.1  17.8  19.8  21.1  21.0 Avg Time 32min  2min  3min  1.3h 48h  11h  14min  2.2h  5h  1h Advanced Auditorium 302 19.6  12.2  16.7  18.7 13.6k 13.7  -  16.6  20.0  21.4  19.8 Ballroom 324 16.3  18.3  13.1  17.9 10.8k 17.2  -  10.4  18.9  18.0  15.6 Courtroom 301 18.2  17.2  12.3  17.1 12.6k 14.6  -  10.2  16.3  18.7  17.8 Palace 509 14.2  11.7  10.8  10.7 21.9k 13.8  -  08.6  11.0  15.3  12.3 Temple 302 18.1  15.7  11.8  09.7 17.5k 13.3  -  11.9  14.8  19.6  16.1 Average 348 17.3  15.0  12.9  14.8 15.6k 14.5  -  11.5  16.2  18.6  16.3 Avg Time 1h  2min  4min  1h 71h  27min  2.8h  10h  2.1h
@@
Next, we consider more than 1k images per scene (right side of Table 4). Here, we run COLMAP with parameters tuned for large images collections (fast) due to the extremely large image sets. ACE0 offers a reconstruction quality comparable to COLMAP on average while also being fast. We run RealityCapture on some of the scenes but it produces fractured reconstructions for these large images sets, leading to low PSNR numbers. DROID-SLAM still struggles on many scenes despite having access to sequential images that are temporally close.
@@ 5 Conclusion and Future Work
We have presented scene coordinate reconstruction, a new approach to learning-based SfM. We learn an implicit, neural scene representation from a set of unposed images. Our method, ACE0, is able to reconstruct a wide variety of scenes. In many cases, the accuracy of estimated poses is close to that of COLMAP and synthesized images visually similar. Unlike previous learning-based SfM methods, ACE0 can be applied to multiple thousand unsorted images, without pose priors, and reconstructs them within a few hours.
Limitations. We show some failure cases in the supplement. Scene coordinate regression struggles with repetitive structures since the network is not able to make multi-modal predictions for visually ambiguous inputs. Scene coordinate regression also struggles with representing large areas. The common solution is to use network ensembles based on pre-clustering of the scene [16, 12] which is difficult in a reconstruction setting. While scene coordinate regression generalises quite well, it has difficulties to bridge extreme view point or lighting changes, such as day versus night. In our experiments, we assumed a simple pinhole camera model with shared intrinsics across images. To the best of our knowledge, scene coordinate regression has not been coupled with image distortion, thus far.
@@ Appendix 0.A Implementation
We start the reconstruction as explained in Section 3.3. We repeat neural mapping and registration of new views until all views have been registered, or less than 1% of additional views are registered compared to the last relocalization stage. In each relocalization iteration, we re-estimate the poses of all views, even if they have been registered before. We also tried a variant that skips estimating poses that have been registered before. But this version achieved slightly lower quality while not giving a significant speed advantage. Re-estimating all poses gives the pipeline the ability to reduce inaccuracies and to correct previous outlier estimates. We consider an image to be registered successfully, if it has at least 500 inlier correspondences with an inlier threshold of 10px. At the end of the reconstruction, we apply a more conservative threshold of 1000 inliers to make sure only high quality poses are kept. For each neural mapping iteration, we load the network weights of the scene coordinate regression network of the previous mapping iteration and refine further. We conclude the reconstruction loop with a final refit of the model, using the standard ACE training parameters. In this last iteration, we start with a randomly initialised network, rather than loading previous weights.
@@
ACE uses a soft-clamped version of the reprojection error in Eq. 3 with a soft threshold that reduces from 50px to 1px throughout training. We use the same, soft-clamped loss but keep the threshold fixed at 50px as we saw no benefit when using the loss schedule but it complicates early stopping. During the seed iteration, we switch from optimizing the Euclidean distance to optimizing the reprojection error if the reprojection error of a scene coordinate is lower than 10px. This hybrid objective is inspired by DSAC* [17].
@@
Runtime Complexity.
The reconstruction time depends on the number of images but also on the spatial distribution of cameras. In the worst case, complexity is O($n^{2}$). E.g. a long camera trajectory without intersections or loops would require proportional to $n$ ACE0 iterations, with $n$ relocalizations in each step, giving O($n^{2}$). The best case complexity is $\Omega(n)$, e.g. for a forward facing scene where all images can be registered in 1-2 ACE0 iterations independent of the number of images. In practise, we observe attractive run times of ACE0.
@@ Appendix 0.B Benchmark
In the following, we give more information about the relocalization experiment of Section 4.1 and 3. We conducted the experiment on the 7-Scenes dataset adhering to its training / test split. Our experiments utilize the pseudo ground truth (pGT) poses of KinectFusion that come with the dataset, as well as the COLMAP pGT of [13]. The authors of [13] obtained the COLMAP pGT by reconstructing the training images of each scene with COLMAP from scratch. To make the pGT scale-metric, they aligned the COLMAP reconstruction with the KinectFusion poses. To obtain poses for the test images of each scene as well, they registered the test images to the training reconstruction. Lastly, they ran a final round of bundle adjustment to refine the test poses while keeping the training poses fixed.
@@
In our experiment, we train the ACE relocalizer on the training images of each scene. Firstly, we train ACE using the KinectFusion poses that come with the dataset. Secondly, we train ACE using the COLMAP pGT of [13]. We evaluate both version of the relocalizer using the COLMAP pGT. As expected, ACE trained with KinectFusion poses scores lower than ACE trained with COLMAP pGT, when evaluating against COLMAP pGT. This confirms the finding of [13] that KinectFusion and COLMAP produce different pose estimates.
@@
Next, we run ACE0 on the training set of each scene to reconstruct it and estimate the training poses. One output of ACE0 is an ACE relocalizer, trained from scratch without poses, in a self-supervised fashion. We apply the ACE0 relocalizer to the test images to estimate their poses - without any further model refinement or joint optimization. The ACE0 reconstruction is scale-less, or only roughly scaled based on the ZoeDepth estimate of the seed image. To enable the scale-metric evaluation of 3, we align the ACE0 pose estimates to the test pseudo ground truth by fitting a similarity transform. Specifically, we run RANSAC [34] with a Kabsch [47] solver on the estimated and the ground truth camera positions with 1000 iterations. Due to the optimization of the similarity transform, we might underestimate the true relocalization error. For a fair comparison, we perform the exact same evaluation procedure for our ACE baselines and fit a similarity transform although they have already been trained with scale-metric poses. The effect of fitting the similarity transform is small. ACE, trained on COLMAP pGT, achieves 97.1% average relocalization accuracy when we fit the similarity transform, and 97.6% average accuracy without fitting the transform. We compute accuracy as the percentage of test images where the pose error is smaller than 5cm and 5 compared to ground truth. The ACE0 relocalizer achieves 93.8% average accuracy when evaluated against the COLMAP pGT. The accuracy differences are smaller than 1% for all scenes except Stairs. This experiment shows that ACE0 is a viable self-supervised relocalizer, and that ACE0 poses are very close to COLMAP poses up to a similarity transform - much closer than e.g. KinectFusion poses are to COLMAP poses.
@@ Appendix 0.C Additional Results
We show estimated poses for more scenes in Fig. 5. We show failure cases in Fig. 6 corresponding to limitations discussed in Section 5.
@@
Additionally, we show synthesized views based on ACE0 poses. Corresponding to our quantitative results (Table 1, 3, Table 4), we select one test image per scene that is representative of the view synthesis quality of ACE0 on that scene. More specifically, we select the test image where ACE0 achieves its median PSNR value. For comparison, we synthesise the same image using the estimated poses of our competitors. For synthesised images of 7-Scenes, see Fig. 7. For comparison with BARF and NoPe-NeRF, we report results on a 200 image subset per scene. We show the associated synthesized images in Fig. 8. For comparison with DUSt3R, we report results on a 50 image subset per scene. We show the associated synthesized images in Fig. 9. We show synthesized images for the MIP-NeRF 360 dataset in Fig. 10. We also show synthesized images for Tanks and Temples, in Fig. 11 for Training scenes, in Fig. 12 for Intermediate scenes, and in Fig. 13 for Advanced scenes.
@@
Figure 7: View Synthesis Quality on 7-Scenes. For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). These results correspond to Table 1. “COLMAP” refers to COLMAP with default parameters.
@@
Figure 8: View Synthesis Quality on 7-Scenes (200 Images Subset). For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). For all methods, we obtain these results on a subset of 200 images per scene to achieve a reasonable training time of NoPe-NeRF. These results correspond to Table 1(right).
@@
Figure 9: View Synthesis Quality on 7-Scenes (50 Images Subset). For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). For all methods, we obtain these results on a subset of 50 images per scene to prevent DUSt3R from running out of GPU memory. These results correspond to Table 1(right).
@@
Figure 10: View Synthesis Quality on the MIP-NeRF 360 dataset. For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). These results correspond to 3.
@@
Figure 11: View Synthesis Quality on Tanks and Temples (Training Scenes). For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). These results correspond to Table 4. Top: Results based on 150-500 images per scenes. Bottom: Results based on 4k-22k frames sampled from the full video.
@@
Figure 12: View Synthesis Quality on Tanks and Temples (Intermediate Scenes). For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). These results correspond to Table 4. Top: Results based on 150-500 images per scenes. Bottom: Results based on 4k-22k frames sampled from the full video.
@@
Figure 13: View Synthesis Quality on Tanks and Temples (Advanced Scenes). For each scene, we show the test image where ACE0 achieves its median PSNR value. For comparison, we show the corresponding results of our baselines as well as the true test image (GT). These results correspond to Table 4. Top: Results based on 150-500 images per scenes. Bottom: Results based on 4k-22k frames sampled from the full video.
@@
We augment Table 1 with SSIM [102] and LPIPS [109] scores for the 7-Scenes dataset, c.f., Tables 5 and 6 respectively. Similarly, Tables 13 and 14 show SSIM [102] and LPIPS [109] scores for the Tanks and Temples dataset, augmenting Table 4. Furthermore, Tables 16 and 16 show SSIM [102] and LPIPS [109] scores for the Mip-NeRF 360 dataset, augmenting 3. In all cases, SSIM and LPIPS behave very similar to PSNR in our experiments.
Pseudo Ground Truth All Frames 200 Frames 50 Frames Frames Kinect COLMAP COLMAP DROID-SLAM ACE0 KF+ACE0 BARF NoPE-NeRF ACE0 DUSt3R ACE0 Fusion (default) (fast) [94] (ours) (ours) [55] [10] (ours) [101] (Ours) Chess 6k 0.69 0.86  0.86  0.76  0.84  0.84 0.55 0.51 0.82 0.66 0.70 Fire 4k 0.55 0.73  0.73  0.44  0.69  0.70 0.45 0.41 0.70 0.55 0.59 Heads 2k 0.69 0.82  0.82  0.78  0.81  0.81 0.54 0.52 0.80 0.69 0.79 Office 10k 0.72 0.85  0.84 failed  0.83  0.83 0.59 0.55 0.80 0.57 0.61 Pumpkin 6k 0.71 0.84  0.84  0.69  0.83  0.82 0.75 0.65 0.84 0.76 0.77 RedKitchen 12k 0.61 0.80  0.80  0.51  0.77  0.78 0.52 0.50 0.72 0.54 0.55 Stairs 3k 0.63 0.59  0.77  0.47  0.57  0.70 0.65 0.66 0.68 0.54 0.55 Average 0.66 0.78  0.81 N/A  0.76  0.78 0.58 0.54 0.76 0.62 0.65 Avg. Time realtime 38h  13h  18min  1h  7min 8.5h 47h 27min 4min* 16min
@@
Pseudo Ground Truth All Frames 200 Frames 50 Frames Frames Kinect COLMAP COLMAP DROID-SLAM ACE0 KF+ACE0 BARF NoPE-NeRF ACE0 DUSt3R ACE0 Fusion (default) (fast) [94] (ours) (ours) [55] [10] (ours) [101] (Ours) Chess 6k 0.37 0.17  0.17  0.33  0.18  0.18 0.80 0.76 0.20 0.42 0.36 Fire 4k 0.44 0.27  0.27  0.70  0.29  0.29 0.79 0.80 0.27 0.40 0.37 Heads 2k 0.42 0.26  0.26  0.32  0.26  0.27 0.75 0.70 0.28 0.40 0.28 Office 10k 0.42 0.25  0.24 failed  0.27  0.26 0.77 0.76 0.31 0.64 0.54 Pumpkin 6k 0.40 0.23  0.24  0.50  0.24  0.25 0.45 0.74 0.22 0.30 0.27 RedKitchen 12k 0.51 0.25  0.25  0.82  0.27  0.27 0.84 0.84 0.35 0.62 0.70 Stairs 3k 0.39 0.50  0.30  0.83  0.51  0.34 0.68 0.70 0.43 0.74 0.61 Average 0.42 0.28  0.25 N/A  0.29  0.26 0.73 0.76 0.29 0.50 0.45 Avg. Time realtime 38h  13h  18min  1h  7min 8.5h 47h 27min 4min* 16min
@@
We show variations of ACE0 in Table 7 alongside PSNR numbers and reconstruction times for 7-Scenes. ACE0 picks 5 random seed images to start the reconstruction. We show average statistics of ACE0 over five runs, starting off with different sets of seed images each time. The standard deviation is low with 0.2 dB in PSNR, and about 4 minutes in terms of reconstruction time.
@@
We run a version of ACE0 that ingests ZoeDepth estimates for all frames and uses them in all reconstruction iterations. Thus, rather than optimizing the reprojection error of Eq. 3, we optimize the Euclidean distance to pseudo ground truth scene coordinates computed from depth estimates. Accuracy is slightly lower, and reconstruction times longer. The depth estimates are not multi-view consistent, and the model has problems to converge.
@@
Train 96.3% 100% 96.4% 100% 97.9%
Advanced Auditorium 98.0% 98.7% 98.7% 99% 98.3%
Ballroom 100% 100% 100% 100% 97.8%
Courtroom 96.3% 100% 93.1% 100% 97.8%
Palace 25.7% 99% 56.0% 74.4% 77.2%
Temple 7.3% 100% 74.4% 100% 92.4%